Dave Carey air conditioning in Gary, Indiana assembles ball bearings for use on “air conditioning” units during manufacturing of “air conditioning” units. Dave Carey wants to locate a low-cost supplier for the ball bearings he uses in his assembly operation. The supplier will be selected based on total annual cost to supply Dave Carey’s needs. Dave Caery’s annual requirements are for 27,500 ball bearings, and the company operates 300 days a year. The following data are available for one of the suppliers being considered. Go to two decimal places in all calculations. SUPPLIER SHIPPING QUANTITY PER SHIPMENT (Q) ANNUAL Freight COSTS PRICE / bearing (p) Annual INVENTORY HOLD COSTS per unit (H) LEAD TIME (DAYS) (L) ADMIN. COSTS Livonia Bearing 2,125 $19,500 $32.50 $7.20 7 $16,500 1A. What is the daily demand (d)? 1B. What are the annual material costs for Livonia Bearing? 1C. What is the annual freight costs? 1D. What are the total annual inventory costs for Livonia Bearing? (Not per unit) 1E. What is the Admin. Costs in dollars? 1F. What are the total annual costs if Livonia Bearing is the supplier selected?
Answer #1

1) The daily demand is calculated by dividing total annual requirement by number of days the company operates= 27500/300 = 91.66 ball bearings.

2) The annual material costs for Livonia Bearings = price per bearing multiplied with total number of bearings = $32.50 * 27500 = $8,93,750.00

3) The annual freight costs = $ 19,500.00 (as per the chart)

4) The annual inventory costs for Livonia Beaarings is $7.20

5) The admin costs is $16.500

6) The total annual costs if Livonia Bearings is the supplier = Annual freight costs + material costs + inventory costs + Admin costs = 19,500+8,93,750+7.20+16500 = $9,29,757.20